# Reawakening knowledge: Anticipatory recovery from catastrophic interference via structured training

### v1

@@ 6 Discussion
In this work, we explored the training dynamics of overparametrized neural networks, especially LLMs, in sequential cyclic fine-tuning, where a finite set of documents are presented in the same order within each epoch. We demonstrated the remarkable phenomenon of anticipatory recovery—networks recover from the initial forgetting before seeing the same document again. The effect holds across many different network instances and training hyper-parameters. This phenomenon is a sharp contrast with the well known phenomenon of catastrophic interference, where forgetting increases monotonically as a network is trained on a sequence of different documents.
@@
We showed that anticipatory recovery occurs only when the network has sufficient width and depth and when it is well fitted to each document before moving to the next. We also discussed the effect of other important factors that influence the degree of recovery, such as the optimizer. Visualizations of model weights, model activations, and gradients exhibit clear temporal structure, which provide insights on the underlying mechanisms of anticipatory recovery. Our computational toy model decouples the shared representations from task-specific parameters and reproduces the phenomenon; however, the mathematical foundation of anticipatory recovery requires further investigation. Additionally, the toy model does not explain why the effect is stronger in LLMs in autoregressive tasks than in other types of architecture or learning objectives.
@@
The cyclic training setup investigated in this work is distinct from the IID training setting assumed in the vast majority of the machine learning literature. It accounts for task repetition and task switching costs, which are critical components of the learning experience of humans and other real world agents. However, our current setup is still highly simplified from real world experience. We have yet to perform experiments on more naturalistic environments with sequential structures. Future research could investigate the emerging training dynamics of neural networks in different types of structured environments, such as multiscale temporal dynamics (Jones et al., 2023), from both theoretical and empirical perspectives.
Designing good learning curricula that achieve both learning efficiency and a low task-switching cost is a promising future direction that can be applied to more naturalistic training setups. On the modeling side, it will be worthwhile to rethink classic continual learning algorithms through the lens of LLMs and exploit and improve their emergent sequential memorization capability.
@@ Acknowledgment
We thank members of the NYU Agentic Learning AI Lab for helpful discussions. The compute was supported by the NYU High Performance Computing resources, services, and staff expertise.
Appendix A Additional Experiment Setups
A.1 Causal Image Modeling Experiments
@@
Image GPT (Chen et al., 2020) is a GPT-2-like model trained to predict the next pixel value in an image. It is pre-trained on the Imagenet dataset (Deng et al., 2009) resized to 32x32. The Image GPT authors provide three pre-trained models of different sizes. In our experiments, we use the Image GPT-small and Image GPT-medium models.
@@
We use the CIFAR-10 (Krizhevsky et al., 2009) dataset for fine-tuning. For tokenization, the pixel RGB values are categorized into 512 pre-determined clusters with the nearest-neighbor classifier, as in Chen et al. (2020). After pre-processing, each image is transformed to a sequence of length 1024, with code book of size 512.
@@
We did not manage to sequentially fine-tune the model stably with the dropout layers, so the dropout layers are turned off during the Image GPT fine-tuning experiments. We use the Adam optimizer (Kingma and Ba, 2015) with learning rate $0.001$. The default hyper-parameters in the experiments are $T=25$ images, $M=10$ gradient update steps, $E=5$ epochs. Same as the LLM experiments, we use the average cross-entropy loss as our evaluation metric.
A.2 Image Classification Experiments
@@
A.3 Computational Toy Model
@@ Appendix B Additional Experiment Results
[Uncaptioned image] Figure 12: Effect of partial document shuffling.
[Uncaptioned image] Figure 13: Effect of learning rate in 1-step GD.
[Uncaptioned image] Figure 14: Effect of number of attention heads.
[Uncaptioned image] Figure 15: Effect of model initialization.
@@ B.2 One-step Gradient Descent with Larger Learning Rate
In Figure 4(b) we observe that there is no anticipation when we take only one gradient descent step on each document with learning rate 0.001. Here we experiment with one-step gradient descent using a higher learning rate, 0.01. We plot the resulting average loss curves under the same training setup in Figure 13. We observe that, with a larger learning rate, slight anticipation is still observed for $1$ gradient step.
@@ B.4 Effect of LLM Model Initialization
Here we compare the performance of the initialization scheme used by Biderman et al. (2023) (also used for all randomly initialized models in the main text) and a simple initialization scheme that samples the weight matrices from an isotropic Gaussian distribution with $\sigma=0.02$. The loss curves for document 1 under these two initializations of the Pythia-1B model are plotted in Figure 15. We observe that Pythia’s initialization scheme achieves much better average loss and also exhibits stronger anticipatory recovery. This demonstrates the importance of LLM initializations. The result is consistent with our observations in section 3.3 that the model’s ability to fit on each task is correlated with the amount of anticipatory recovery.
@@
B.5 Effect of Number of Gradient Steps with Inverse Learning Rate Scaling
In Figure 15(a) we experiment with inversely scaling the learning rate with the number of gradient steps. We use a learning rate of $0.01$ for $M=1$, learning rate $0.002$ for $M=5$, learning rate $0.001$ for $M=10$, and learning rate $0.0005$ for $M=20$. The results suggest that the anticipation effect is stronger when the same total update is divided among more gradient steps.
B.6 Effect of Number of Gradient Steps for Long Context Length
In Figure 15(b) we experiment with different number of gradient steps $M\in\{10,20,40\}$ for context length 1024. The results confirm that longer context length is not a fundamental limitation to anticipatory recovery, and we can achieve the same recovery score as a smaller context length with more gradient steps.
Appendix C Additional Visualizations
@@
Figure 17: Magnitude of (a) model activation updates and (b) model weight updates through four epochs of cyclic training.
@@ C.2 Pairwise Recovery Matrices for Longer Document Sequences
Similar to Figure 7(b), we plot the pairwise loss recoveries for each pair of documents $(\bm{x}_{i},\bm{x}_{j})$ in longer document sequences, where $T=50,100,200$ respectively, in Figure 18. We use the 1B model and default hyperparameters. We observe that, as we increase the length of the document sequence, the highlight area near the center of the matrix is separated into two blobs, one in the top-left corner and the other in the bottom-right corner. We also observe a ”boundary” on the sides of the matrix where there is little or no recovery. The width of this ”boundary” stays relatively constant across different lengths of document sequences and is around 10 to 15 documents. This confirms our observation in the main text that the recovery on document $\bm{x}_{j}$ when fine-tuning on a proximal document $\bm{x}_{i}$ is highest when the model checkpoint is taken from document $\bm{x}_{j\pm b}$ where $b$ is a small number relative to the length of the document sequence.

### v2

@@ 6 Discussion and Limitations
In this work, we explored the training dynamics of overparametrized neural networks, especially LLMs, in sequential cyclic fine-tuning, where a finite set of documents are presented in the same order within each epoch. We demonstrated the remarkable phenomenon of anticipatory recovery—networks recover from the initial forgetting before seeing the same document again. The effect holds across many different network instances and training hyperparameters. This phenomenon is a sharp contrast with the well known phenomenon of catastrophic interference, where forgetting increases monotonically as a network is trained on a sequence of different documents.
@@
We showed that anticipatory recovery occurs only when the network has sufficient width and depth and when it is well fitted to each document before moving to the next. Visualizations of model weights, model activations, and gradients exhibit clear temporal structure, which provide insights on the underlying mechanisms of anticipatory recovery.
Our research indicates that there is value in exploring naturalistic task sequences within continual learning, where tasks interleave in statistically regular patterns. This approach could expand the field’s current focus on learning and retaining new tasks to also consider how effectively previously encountered tasks are re-learned when they reappear. With the anticipatory recovery phenomenon, we discovered a mechanism in which ML models can do surprisingly better than expected on prequential evaluation. By analyzing the different factors of model pre-training and fine-tuning that moderate this phenomenon, our experiments provide a promising first step toward leveraging structured training with agents in realistic environments.
Limitations.
@@
The cyclic training setup investigated in this work is distinct from the IID training setting assumed in the vast majority of the machine learning literature. It accounts for task repetition and task switching costs, which are critical components of the learning experience of humans and other real world agents. However, our current setup is still highly simplified. Future research could investigate the emerging training dynamics of neural networks in different types of structured environments, such as multiscale temporal dynamics (jones2023learning, ), from both theoretical and empirical perspectives.
The mathematical foundation of anticipatory recovery also requires further investigation. Although our computational toy model reproduces the anticipatory recovery phenomenon, it does not explain why the effect is stronger in LLMs and autoregressive tasks than in other types of architecture or learning objectives. In terms of the empirical experiments, the ablation studies are only run in a single setting. Future research could run the experiments in more settings to reach more conclusive results and investigate possible alternative theoretical explanations to the anticipatory recovery phenomenon.
@@ Acknowledgment
We thank the Microsoft Accelerating Foundation Models Research program for providing Azure cloud compute credits. We thank members of the NYU Agentic Learning AI Lab for helpful discussions. The compute was supported by the NYU High Performance Computing resources, services, and staff expertise. MJ was supported by NSF grant 2020-906.
Appendix A Additional Experiment Details
A.1 LLM Experiments
We use the Huggingface Transformers Library [63] for fine-tuning the LLMs. The learning rate $0.001$ for vanilla gradient descent and $0.00001$ for Adam. For all experiments we run 3 to 5 trials with different random seeds, except the results in Figure 1 which are based on 20 seeds. The shaded area in the figures denotes standard deviation among trials and documents (for shift-averaged loss curves).
A.2 Causal Image Modeling Experiments
@@
Image GPT [14] is a GPT-2-like model trained to predict the next pixel value in an image. It is pre-trained on the Imagenet dataset [19] resized to 32x32. The Image GPT authors provide three pre-trained models of different sizes. In our experiments, we use the Image GPT-small and Image GPT-medium models.
@@
We use the CIFAR-10 [18] dataset for fine-tuning. For tokenization, the pixel RGB values are categorized into 512 pre-determined clusters with the nearest-neighbor classifier, as in [14]. After pre-processing, each image is transformed to a sequence of length 1024, with code book of size 512.
@@
We did not manage to sequentially fine-tune the model stably with the dropout layers, so the dropout layers are turned off during the Image GPT fine-tuning experiments. We use the Adam optimizer [20] with learning rate $0.001$. The default hyperparameters in the experiments are $T=25$ images, $M=10$ gradient update steps, $E=5$ epochs. Same as the LLM experiments, we use the average cross-entropy loss as our evaluation metric.
A.3 Image Classification Experiments
@@
A.4 Shift-averaged Loss Calculation
The shift-averaged loss curves plotted in Figure 1(b) are calculated by replicating Figure 1(a) on each document in the training sequence, re-aligning these curves so that 0 on the x-axis always represents the moment before the first occurrence of the focal document, and average them. For example, if the length of the sequence is 50, then for training epoch 0.5 on the x-axis, we take the loss of document 1 after training on document 25; the loss of document 2 after training on document 26; …; the loss of document 50 after training on document 24 of the next epoch; and average these losses. Subsequent figures (Figure 2-712-19) are plotted with the same approach.
A.5 Computational Toy Model
@@
A.6 Compute Resources
Each experiment presented in the paper is run with one NVIDIA A100 GPU, 2 CPUs, and 32GB of RAM. The training time highly depends on the hyperparameter choices, especially model size and number of gradient steps. The longest fine-tuning experiment with 20 gradient steps per episode on the Pythia-1B model takes roughly 30 minutes under this setup. The minimal compute resource requirement needed to reproduce the experiments with a Pythia-1B model is one GPU with 16GB of GPU memory.
@@ Appendix B Additional Experiment Results
Figure 12: Effect of partial document shuffling.
Figure 13: Effect of learning rate in 1-step GD.
Figure 14: Effect of number of attention heads.
Figure 15: Effect of model initialization.
Figure 16: (Left) Effect of pre-training steps. The full pre-training process is 143k steps. (Right) Recovery scores for models with different pre-training steps.
Figure 17: Loss curve for cosine learning rate schedule.
Figure 18: Effect of inserting random documents after the repeating sequence.
@@ B.2 One-step Gradient Descent with Larger Learning Rate
In Figure 5(b) we observe that there is no anticipation when we take only one gradient descent step on each document with learning rate 0.001. We experimented with one-step gradient descent using a higher learning rate, 0.01. We plot the resulting average loss curves under the same training setup in Figure 13. We observe that, with a larger learning rate, slight anticipation is still observed for $1$ gradient step.
@@ B.4 Effect of LLM Model Initialization
Here we compare the performance of the initialization scheme used by [11] (also used for all randomly initialized models in the main text) and a simple initialization scheme that samples the weight matrices from an isotropic Gaussian distribution with $\sigma=0.02$. The loss curves for document 1 under these two initializations of the Pythia-1B model are plotted in Figure 15. We observe that Pythia’s initialization scheme achieves much better average loss and also exhibits stronger anticipatory recovery. This demonstrates the importance of LLM initializations. The result is consistent with our observations in section 3.3 that the model’s ability to fit on each task is correlated with the amount of anticipatory recovery.
B.5 Effect of Pre-training Steps
In addition to comparing pre-trained models and randomly initialized models in Figure 2, we further study the effect of model pre-training by examining model checkpoints with different numbers of pre-training steps. We took Pythia models pre-trained for 6K, 12K, 24K, 48K, and 96K steps and plot the shift-averaged loss curves for cyclic fine-tuning in Figure B. We found that more pre-training does give rise to higher anticipatory recovery. As we summarize at the end of section 3.3, we hypothesize this result fits a broader pattern in which the strength of the anticipatory recovery effect is related to how well the model can fit each successive training task. Models with more pre-training steps are more capable of fitting each successive training task, and therefore exhibit higher anticipatory recovery.
B.6 Effect of Cosine Learning Rate Schedule
For experiments in the main paper we used a constant learning rate during the fine-tuning process. To study whether anticipatory recovery occurs in typical LLM optimization schemes, we experimented with cosine learning rate scheduling on the Pythia-1B model with 10 epochs (minimum learning rate = 0, maximum number of epochs = 10), and plot the results in Figure 17. We show that the model also exhibits the anticipatory recovery effect in other learning rate schedules.
B.7 Effect of Inserting Random Documents after the Repeating Sequence
To examine how the anticipatory recovery effect may generalize to other forms of structured training, we experimented with a new setting where only the first 20 documents are kept fixed in each epoch, and a random number of other documents (between 20 and 100) are inserted after the first 20. These random padding documents appear only once in the entire sequence. This new setting generalizes cyclic training in that (1) rather than having the same documents in every epoch, we insert random other documents between every repetition (2) epochs can have different lengths. The resulting loss curve is plotted in Figure 18. We still observe anticipatory recovery for documents 2 through 20 in this setting, suggesting that anticipatory recovery exists as long as there is a repeating sub-sequence in the data stream.
@@ B.7 Effect of Inserting Random Documents after the Repeating Sequence
B.8 Effect of Number of Gradient Steps with Inverse Learning Rate Scaling
In Figure 19(a) we experimented with inversely scaling the learning rate with the number of gradient steps. We use a learning rate of $0.01$ for $M=1$, learning rate $0.002$ for $M=5$, learning rate $0.001$ for $M=10$, and learning rate $0.0005$ for $M=20$. The results suggest that the anticipation effect is stronger when the same total update is divided among more gradient steps.
B.9 Effect of Number of Gradient Steps for Long Context Length
In Figure 19(b) we experimented with different number of gradient steps $M\in\{10,20,40\}$ for context length 1024. The results confirm that longer context length is not a fundamental limitation to anticipatory recovery, and we can achieve the same recovery score as a smaller context length with more gradient steps.
Figure 20: Experiments with GPT2-large.
Figure 21: Experiments with wikitext-103 dataset.
@@ B.9 Effect of Number of Gradient Steps for Long Context Length
Figure 22: Magnitude of (a) model activation updates and (b) model weight updates through cyclic training.
B.10 Experiments with GPT-2
To evaluate how the anticipatory recovery phenomenon generalizes across different LLM architectures, we experimented with GPT-2 architecture [64], specifically the GPT2-large pre-trained model (812M parameters) on the CNN/Daily Mail dataset. The loss curve for document 1 is plotted in Figure 20. The model consistently observed anticipatory recovery. Note that GPT-2 is the predecessor of many modern LLMs and therefore the results further suggest that the anticipatory recovery phenomenon is prevalent among more recent LLM architectures.
B.11 Experiments with Wikitext
To evaluate how the anticipatory recovery phenomenon generalizes across different natural language datasets, we experimented with the wikitext-103 dataset [65], which contains over 100 million tokens from articles on Wikipedia. Since Wikipedia data is part of the pre-training dataset of Pythia, we only experiment with randomly initialized models. The loss curve for document 1 is plotted in Figure 21. The result suggest that the anticipatory recovery phenomenon is generalizable to different data sources.
Appendix C Additional Visualizations
@@ C.2 Pairwise Recovery Matrices for Longer Document Sequences
Similar to Figure 8(b), we plot the pairwise loss recoveries for each pair of documents $(\bm{x}_{i},\bm{x}_{j})$ in longer document sequences, where $T=50,100,200$ respectively, in Figure 23. We use the 1B model and default hyperparameters. We observe that, as we increase the length of the document sequence, the highlight area near the center of the matrix is separated into two blobs, one in the top-left corner and the other in the bottom-right corner. We also observe a "boundary" on the sides of the matrix where there is little or no recovery. The width of this "boundary" stays relatively constant across different lengths of document sequences and is around 10 to 15 documents. This confirms our observation in the main text that the recovery on document $\bm{x}_{j}$ when fine-tuning on a proximal document $\bm{x}_{i}$ is highest when the model checkpoint is taken from document $\bm{x}_{j\pm b}$ where $b$ is a small number relative to the length of the document sequence.
Appendix D Additional Related Work
Learning in Structured Environments.
Our research also relates to the more general topic of learning in structured environments. [62] studied regression and classification tasks with multi-scale temporal structure in the environment characterized by $1/f$ dynamics. While the cyclic training setting that we study is a more simplified setup than that of [62], we aim at unveiling more insights on applying standard SGD on over-parameterized networks. A potential direction for future work would be to study anticipatory recovery in regimes with richer, hierarchical sequence structure.
LLM Emergent Capabilities.
Recent advancements in large-scale Transformer networks [66, 1] have demonstrated exceptional ability to model long sequence language data. Beyond basic language modeling and downstream task performance, these models have shown emergent behaviors [67] that appear to manifest only beyond a certain model scale [2, 68, 69, 70]. Related to our research, recent studies reveal that LLMs possess remarkable memorization skills, enabling them to recall news sentences after just a few exposures [11, 71, 72]. However, the sequential learning dynamics behind such memorization have not been thoroughly examined. Our work comprehensively explore the sequential learning setting with cyclic task repetition and demonstrates task anticipation, a new emergent capability of large models.
Appendix E Broader Impact
This research deepens our understanding of catastrophic interference in naturalistic training setups. This understanding could lead to the design of better training algorithms of LLMs and other large neural networks that are more similar to human learning. These algorithms may give rise to more powerful and embodied AI systems with online adaptive learning capability, which may have many potential societal consequences.
Recovery from catastrophic interference might be undesirable in scenarios where some documents are intended to be forgotten. Our research provide more understanding into the anticipatory recovery phenomenon. Future research into the unlearning problem as well as a deeper understanding of the training dynamics of different types of structured environments can help further address the privacy issue.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The authors confirm that the claims made in the abstract and introduction accurately reflect the paper’s contributions and scope.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations of this work are discussed in Section 6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include theoretical results. While we provide a well-defined computational model in section 4.4, we are not trying to prove new theoretical results and the purpose of section 4.4 is to provide more intuition on how the anticipatory recovery phenomenon might occur.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: In Sections 23 and Appendices AB we disclose all the information needed to reproduce the main experimental results of the paper. The pre-trained models and datasets we use are also publicly available. We also include the code for reproducing main experimental results in the supplementary material.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide the code and instructions for reproducing main experimental results in the supplementary material.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The experimental settings and details are provided in Section 2 and Appendix A.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: All experiment result figures and tables in Section 3 of the paper are accompanied by error bars or confidence intervals. Some error bars might not be visible since they are too small.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provided detailed information on the compute resources needed to reproduce the experiments in Section A.6.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The authors confirm that this paper conforms with the NeurIPS Code of Ethics. We used publicly available, non-deprecated datasets and models, and conform with their licenses.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The broader impacts of this work are discussed in Appendix E.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper does not release new data or models.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The original papers for all the code, data and models used in the main experiments of the paper are cited. All these existing assets involved are publicly available, under the Apache 2.0 License (Pythia, huggingface transformers) or MIT License (cnn/dailymail version 3.0.0).
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: This paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.